# Fast and accurate parameter estimation of high-redshift sources with the Einstein Telescope

### v1

@@ III.2 Population analysis
Figure 6: Probability-probability (P-P) plot showing the confidence interval (C.I.) on the $x$-axis against the fraction of events within that C.I on the $y$-axis, for posterior distributions obtained with Dingo (without importance sampling) from 1000 injected binary black holes (BBHs) signals sampled from the prior (Tab. 1). The shaded regions represent the $1\sigma$, $2\sigma$, and $3\sigma$ C.I., with $p$-values given for each parameter in parenthesis. See Sec. III.2 for details.
@@
We generate 1000 simulated BBH signals with parameters randomly sampled from the prior distribution (Tab. 1) and process them using Dingo-IS. For each injection, we draw $10^{4}$ samples, estimate the phase $\phi_{c}$ in post-processing, and compute the importance weights, as detailed in Sec. II.4.
Fig. 6 presents a probability-probability (P-P) plot, which is a standard method to validate inference algorithms (Samantha R Cook and Rubin, 2006; Veitch et al., 2015; Talts et al., 2020; Green et al., 2020). The P-P plot checks whether the fraction of injected values recovered within a given confidence interval follows a uniform distribution. This also allows us to compute the $p$-values for each parameter. For an ideal sampler, an $n\%$ confidence interval should contain $n\%$ of the events on average, making a diagonal line in the plot. Statistical fluctuations cause deviations, so we also display the $1\sigma$, $2\sigma$, and $3\sigma$ confidence intervals.
Fig. 6 confirms that the Dingo proposal distribution $q(\theta|d)$, trained for ET-$\Delta$, performs reasonably well for most parameters, with the curves following the diagonal. However, some deviations are visible, most notably for the phase $\phi_{c}$, as also indicated by its respective $p$-value. Nevertheless, any deviations in samples drawn with NPE are corrected through importance sampling.
Fig. 7 shows how the sample efficiency varies with detector-frame chirp mass and luminosity distance of the processed injections. Dingo-IS enables PE in a wide redshift range ($1\lesssim z\lesssim 45$) with high accuracy. The average and median sample efficiency are 12.5% and 10%, respectively, with $\sim 85$% of sources with sample efficiency $>1\%$. The highest sample efficiencies (about 10-20%) are observed for the most distant BBHs. They have low SNR, making them easier for NPE to learn, as their posteriors are closer to the prior. This can be also seen as the average sample efficiency increases to $\sim 18\%$ $(\sim 20\%)$ if we consider only sources merging at $z>4$ $(z>10)$. Similar sample efficiencies were reported in Dax et al. (2023), where they analyzed 42 BBH events detected with the LIGO and Virgo interferometers (Abbott et al., 2021, 2023).
Fig. 7 also highlights regions of low sample efficiency, particularly at $z\lesssim 2$. This behavior might be influenced by the chosen prior on the luminosity distance. As shown in previous studies (e.g., Green and Gair, 2021), using a uniform prior in luminosity distance might improve performance by exposing the neural network to more low-$d_{\mathrm{L}}$ events during training. In regions with low sample efficiency, we can compensate by increasing the number of posterior samples drawn with Dingo-IS.
@@
Figure 8: The top, middle, and bottom panels show the relative errors for the detector-frame chirp mass ($\mathcal{M}_{\mathrm{d}}$), luminosity distance ($d_{\mathrm{L}}$), and sky localization ($\Delta\Omega_{90\%}$), respectively. The $x$-axis corresponds to results from Fisher-information matrix approximation (FIM, GWFish+Priors), while the $y$-axis represents neural posterior estimate with importance sampling (NPE, Dingo-IS). The points are color-coded by the injected luminosity distance. Event 1 (Event 2) discussed in Sec. III.1 (Appendix B) is marked with a red (orange) circle. See Sec. III.2 for details.
@@
Figure 8 presents scatter plots that compare the relative errors in chirp mass, luminosity distance, and sky localization estimated with NPE and FIM. In particular, we compute the 90% highest probability density sky area by evaluating a kernel density estimate on a multi-order HEALPix grid (Gorski et al., 1999), using the ligo-skymap library (Singer and Price, 2016; Singer et al., 2016a, b). The other reported intervals, such as $\Delta x/\langle x\rangle$, represent the relative variation of the parameter around its average, where $\Delta x$ is the square root of the corresponding covariance element, reflecting the $1\sigma$ uncertainty.
@@
For the chirp mass, FIM slightly underestimates the error compared to NPE. In the worst cases, the difference remains within a factor of 4 and is independent of the injected luminosity distance.
@@
The relative errors on the luminosity distance obtained with both methods agree well for injected distances up to $d^{\mathrm{inj}}_{\mathrm{L}}\sim 10^{5}$ Mpc. However, for $d^{\mathrm{inj}}_{\mathrm{L}}>10^{5}$ Mpc, the estimated relative error from GWFish+Priors increases systematically, reaching $\sim 1$. This indicates that the FIM analysis returns the chosen prior range (Tab. 1). The relative error on luminosity distance is between 2 and 4 times larger compared to what is derived with NPE. Therefore, FIM tends to overestimate luminosity distance errors for high-redshift sources. The accurate evaluation of distance errors is important for analyses of the ET science case, as discussed in Sec. IV.4.
@@
For sky localization, the discrepancy is the most severe. As seen in Fig. 5, ET-$\Delta$ sky localization for short-lived sources is characterized by eight degenerate sky modes (see Sec. IV.1 for a discussion), which FIM fails to capture by construction. For well-localized sources ($\Delta\Omega_{90\%}<100~{}\mathrm{deg}^{2}$), the discrepancy appears to be very close to a factor of 8. For poorly localized sources ($\Delta\Omega_{90\%}>100~{}\mathrm{deg}^{2}$), the error difference can increase by up to three orders of magnitude, worsening with the injected distance. This has also important implications for the ET science case, discussed in Sec. IV.4.
@@
Fig. 9 shows that the FIM analysis accurately estimates the error for the mass ratio but underestimates the errors for the aligned spins. In particular, the uncertainty on $\chi_{1}$ and $\chi_{2}$ for most sources is $\Delta\chi_{i}\sim 0.3$, which is the spread of the prior distribution (Eq. 8). This indicates that aligned spins estimated with NPE are mostly unconstrained for these sources. These findings are consistent with previous studies, such as Dupletsa et al. (2025).
@@
Figs. 8 and 9 display 344 of the 500 selected events that have a sample efficiency above 10%. The remaining events are excluded because their covariance matrices, given by the inverse of the FIM in Eq. 16, are close to singular. This exclusion ensures numerical stability when sampling from a truncated multivariate Gaussian, as described in Dupletsa et al. (2025).
@@
Moreover, the results shown in Figs. 8 and 9 are based on aligned spins, which means we do not account for spin induced precession of the orbital plane. If measured, precession could help break degeneracies between waveform parameters. We discuss this further in Sec. IV.2.
@@ IV.1 Sky modes
Figure 10: Right ascension and declination samples from Dingo-IS converted in azimuth ($\lambda$) and altitude ($\beta$). Degeneracies, as listed in Tab. 3, are included. Color-coded is the square root of the sum of the antenna power pattern functions (Eq. IV.1) of ET-1 (top panel), ET-2 (middle panel) and ET-3 (bottom panel). See Sec. IV.1 for details.
@@
Reflected $-\beta,\lambda$
Antipodal $-\beta,\lambda+\pi$
@@
6th $-\beta,\lambda+\pi/2$
7th $-\beta,\lambda-\pi/2$
Table 3: Sky mode degeneracies reported in Fig. 10. We use the same nomenclature as in Marsat et al. (2021). Coordinates are given in the horizontal reference frame, defined by altitude ($\beta$) and azimuth ($\lambda$) angles. See Sec. IV.1 for details.
The degeneracies in sky localization, which are evident in the two-dimensional marginal distribution of ra and dec in Fig. 5, reveal eight distinct sky modes. These modes have been previously discussed for ET in its triangular configuration (see Fig. 2 in Singh and Bulik 2022 and Fig. 2 in Singh and Bulik 2021) and in the LISA triangular space-based interferometer, assuming frequency- and time-independent antenna patterns (Baibhav et al., 2020; Marsat et al., 2021). For consistency, we label these modes in Tab. 3 using the same nomenclature as in Marsat et al. (2021).
These sky modes result from the symmetries in the antenna pattern functions of the three detectors that form ET-$\Delta$. The antenna response functions, which take into account that GW detectors are not equally sensitive to gravitational waves from every direction, can be expressed in the horizontal reference frame, defined by the azimuth ($\beta$) and altitude ($\lambda$) angles. In this frame, the antenna amplitude pattern functions are given by (Regimbau et al., 2012):
@@
All these antenna patterns are symmetric under the following transformations: reflection of the sky position across the horizontal plane ($\beta\to-\beta$), rotation about the vertical direction by half a turn ($\lambda\to\lambda+\pi$), and simultaneous rotation of the sky position and polarization angle by a quarter of a turn ($\lambda\to\lambda+\pi/2$, $\psi\to\psi+\pi/2$) (Marsat et al., 2021). Because these transformations produce identical signals in each ET component, they lead to eight degenerate sky positions in the posterior, evenly spaced in $\lambda$ and symmetrically distributed above and below the horizon.
To visualize this effect, we show the total response of each ET component to an incident wave, also called antenna power pattern function: for ET-1, this is given by (Schutz, 2011):
@@
where the dependency on the polarization angle $\psi$ cancels out. Fig. 10 displays the sky location samples in the horizontal reference frame, color-coded by Eq. IV.1. The antenna responses remain identical across all eight modes in all three ET detectors.
The same degeneracies also occur when analyzing the PE of massive black hole binaries (MBHBs) with LISA (Marsat et al., 2021; Pitte et al., 2023; Sun et al., 2025; Yi et al., 2025), provided that the antenna pattern does not change due to Earth’s rotation and the long-wavelength approximation is assumed (Essick et al., 2017). However, in LISA, these approximations can be relaxed, allowing for resolution of these degeneracies (Buscicchio et al., 2021). Specifically, an MBHB with a total mass greater than $10^{6}~{}\mathrm{M}_{\odot}$ can stay in the LISA frequency band for more than 40 hours (Marsat et al., 2021), meaning the rotation of the satellite constellation can have a significant effect.
The long-wavelength approximation (also known as the low-frequency limit) assumes that the antenna pattern functions (Eq. 18 and Eq. 19) act purely as projections of the GW signal onto the detector. This approximation is valid when the length of the detector arms is much shorter than the wavelength of the incoming GW (Misner et al., 1973; Hawking and Israel, 1989; Rakhmanov et al., 2008). In reality, this approximation can break down, and the response can also depend on the GW frequency. The frequency at which the long-wavelength approximation becomes important is determined by the unperturbed round-trip travel time along one arm, also known as the free spectral range (FSR; Marsat et al., 2021; Essick et al., 2017):
@@
Here, $c$ is the speed of light, and for LISA, $L=2.5\times 10^{9}~{}\mathrm{m}$, which gives $f_{\mathrm{FSR}}=0.06$ Hz. The merger frequency of MBHBs within the LISA frequency band is of the same order of magnitude as $f_{\mathrm{FSR}}$, breaking the long-wavelength approximation (Marsat et al., 2021).
In comparison, for ET-$\Delta$, the free spectral range is $f_{\mathrm{FSR}}=1.5\times 10^{4}~{}\mathrm{Hz}$, which is at least three orders of magnitude higher than the typical merger frequency of the sources considered in this study (see Eq. 2). Moreover, BBH mergers with high detector-frame chirp mass last only a few seconds in the ET-$\Delta$ frequency band. Therefore, we conclude that assuming a time and frequency independent antenna pattern is still a valid approximation for high-redshift sources detected with ET-$\Delta$.
IV.2 Higher-order modes and precessing spins
Figure 11: Marginalized one- and two- dimensional posterior distributions for Event 1 with zero noise. We compare waveforms that include aligned spins and higher-order modes (HOM A, dark green) with those that include only the dominant mode ($(2,|2|)$ A, dark magenta), both analyzed using Bilby. Vertical and horizontal lines indicate the true injected values (see Tab. 2), while the contours represent the 68% and 95% credible regions. Further details are provided in Sec. IV.2.
In the time domain, a GW signal $h=h_{+}+ih_{\times}$ can be expressed as (Newman and Penrose, 1966; Berti et al., 2007; Kidder, 2008):

### v2

@@ III.2 Population analysis
Figure 6: Probability–probability (P–P) plot showing the confidence interval (C.I.) on the $x$-axis against the fraction of events within that C.I on the $y$-axis, for posterior distributions obtained with Dingo (without importance sampling) from 1000 injected binary black holes (BBHs) signals sampled from the prior (Table 1). The shaded regions represent the $1\sigma$, $2\sigma$, and $3\sigma$ C.I., with $p$-values given for each parameter in parenthesis and a combined $p$-value of 0.001. See Sec. III.2 for details.
@@
We generate 1000 simulated BBH signals with parameters randomly sampled from the prior distribution (Table 1) and process them using Dingo-IS. For each injection, we draw $10^{4}$ samples, estimate the phase $\phi_{c}$ in post-processing, and compute the importance weights, as detailed in Sec. II.4.
Fig. 6 presents a probability–probability (P–P) plot, which is a standard method to validate inference algorithms (Cook01092006; Veitch:2014wba; talts2020validatingbayesianinferencealgorithms; 2020PhRvD.102j4057G). The P-P plot checks whether the fraction of injected values recovered within a given confidence interval follows a uniform distribution. This also allows us to compute the $p$-values for each parameter. For an ideal sampler, an $n\%$ confidence interval should contain $n\%$ of the events on average, making a diagonal line in the plot. Statistical fluctuations cause deviations, so we also display the $1\sigma$, $2\sigma$, and $3\sigma$ confidence intervals.
Figure 6 confirms that the Dingo proposal distribution $q(\theta|d)$, trained for ET-$\Delta$, performs reasonably well for most parameters, with the curves following the diagonal. However, some deviations are visible, most notably for the phase $\phi_{c}$, as also indicated by its respective $p$-value. This makes importance sampling necessary to correct deviations in samples drawn with NPE.
Fig. 7 shows how the sample efficiency varies with detector-frame chirp mass and luminosity distance of the processed injections. Dingo-IS enables PE in a wide redshift range ($1\lesssim z\lesssim 45$) with high accuracy. The average and median sample efficiency are 12.5% and 10%, respectively, with $\sim 85$% of sources with sample efficiency $1\%$. The highest sample efficiencies (about 10%–20%) are observed for the most distant BBHs. They have low SNR, making them easier for NPE to learn, as their posteriors are closer to the prior. This can be also seen as the average sample efficiency increases to $\sim 18\%$ $(\sim 20\%)$ if we consider only sources merging at $z>4$ $(z>10)$. Similar sample efficiencies were reported in 2023PhRvL.130q1403D, where they analyzed 42 BBH events detected with the LIGO and Virgo interferometers (LIGOScientific:2020ibl; KAGRA:2021vkt).
Fig. 7 also highlights regions of low sample efficiency, particularly at $z\lesssim 2$. This behavior might be influenced by the chosen prior on the luminosity distance. As shown in previous studies (e.g., Green:2020dnx), using a uniform prior in luminosity distance might improve performance by exposing the neural network to more low-$d_{\mathrm{L}}$ events during training. In regions with low sample efficiency, we can compensate by increasing the number of posterior samples drawn with Dingo-IS.
@@
Figure 8: The top, middle, and bottom panels show the relative variations for the detector-frame chirp mass ($\mathcal{M}_{\mathrm{d}}$), luminosity distance ($d_{\mathrm{L}}$), and sky localization ($\Delta\Omega_{90\%}$), respectively. The $x$-axis corresponds to results from Fisher-information matrix approximation (FIM, GWFish+Priors), while the $y$-axis represents neural posterior estimate with importance sampling (NPE, Dingo-IS). The points are color-coded by the injected luminosity distance. Event 1 (Event 2) discussed in Sec. III.1 (Appendix B) is marked with a red (orange) circle. See Sec. III.2 for details.
@@
Figure 8 presents scatter plots that compare the relative variations in chirp mass, luminosity distance, and sky localization estimated with NPE and FIM. In particular, we compute the 90% highest probability density sky area by evaluating a kernel density estimate on a multiorder HEALPix grid (Gorski:1999rt), using the ligo-skymap library (Singer:2015ema; Singer:2016eax; Singer:2016erz). The other reported intervals, such as $\Delta x/\langle x\rangle$, represent the relative variation of the parameter around its average, where $\Delta x$ is the square root of the corresponding covariance element, reflecting the $1\sigma$ uncertainty.
@@
For the chirp mass, FIM systematically underestimates the error compared to NPE. In the worst cases, the difference remains within a factor of 4 and is independent of the injected luminosity distance.
@@
For sources with injected luminosity distance up to $d^{\mathrm{inj}}_{\mathrm{L}}\sim 10^{5}~\mathrm{Mpc}$, the FIM overestimates and underestimates the relative variations compared to NPE by up to a factor of 4. Moreover, for $d^{\mathrm{inj}}_{\mathrm{L}}>10^{5}$ Mpc, the estimated relative variation from GWFish+Priors increases systematically, reaching $\sim 1$. This indicates that the FIM analysis returns the chosen prior range (Table 1). The relative variation on luminosity distance is on average between 2 and 4 times larger compared to what is derived with NPE. Therefore, FIM tends to overestimate luminosity distance errors for high-redshift sources. The accurate evaluation of distance errors is important for analyses of the ET science case, as discussed in Sec. IV.5.
@@
For sky localization, the discrepancy is the most severe. As seen in Fig. 5, ET-$\Delta$ sky localization for short-lived sources is characterized by eight degenerate sky modes (see Sec. IV.1 for a discussion), which FIM fails to capture by construction. For well-localized sources ($\Delta\Omega_{90\%}<100~\mathrm{deg}^{2}$), the discrepancy appears to be very close to a factor of 8. For poorly localized sources ($\Delta\Omega_{90\%}>100~\mathrm{deg}^{2}$), the error difference can increase by up to 3 orders of magnitude, worsening with the injected distance. This has also important implications for the ET science case, discussed in Sec. IV.5.
@@
Figure 9 shows that the FIM analysis tends to underestimate the error for the mass ratio (within a factor of 4) and systematically underestimates the errors for the aligned spins. In particular, the uncertainty on $\chi_{1}$ and $\chi_{2}$ for most sources is $\Delta\chi_{i}\sim 0.3$, which is the spread of the prior distribution (Eq. 8). This indicates that aligned spins estimated with NPE are mostly unconstrained for these sources. These findings are consistent with previous studies, such as Dupletsa:2024gfl.
@@
Figures 8 and 9 display 344 of the 500 selected events that have a sample efficiency above 10%. The remaining 156 events are excluded because their covariance matrices, given by the inverse of the FIM in Eq. 16, are close to singular. This exclusion is required in GWFish+Priors to ensure numerical stability when sampling from a truncated multivariate Gaussian, as described in Dupletsa:2024gfl.
@@
Moreover, the results shown in Figs. 8 and 9 are based on quasi-circular systems with aligned spins, which means we do not account for spin-induced precession of the orbital plane or orbital eccentricity. If measured, precession and/or eccentricity could help break degeneracies between waveform parameters, leading to even more well-constrained posteriors (Lower:2018seu; Yang:2022fgp; Yang:2022tig; Ng:2022vbz; Saini:2023wdk). We discuss this further in Sec. IV.3.
@@ IV.1 Sky modes
Figure 10: Right ascension and declination samples from Dingo-IS converted in azimuth ($\lambda$) and altitude ($\beta$). Degeneracies, as listed in Table 3, are included. Color-coded is the square root of the sum of the antenna power pattern functions (Eq. IV.1) of ET-1 (top panel), ET-2 (middle panel) and ET-3 (bottom panel). See Sec. IV.1 for details.
@@
Reflected $\beta,\lambda$
Antipodal $\beta,\lambda+\pi$
@@
6th $\beta,\lambda+\pi/2$
7th $\beta,\lambda-\pi/2$
Table 3: Sky mode degeneracies reported in Fig. 10. We use the same nomenclature as in Marsat:2020rtl. Coordinates are given in the horizontal reference frame, defined by altitude ($\beta$) and azimuth ($\lambda$) angles. See Sec. IV.1 for details.
The degeneracies in sky localization, which are evident in the two-dimensional marginal distribution of ra and dec in Fig. 5, reveal eight distinct sky modes. These modes have been previously discussed for ET in its triangular configuration (see Fig. 2 in Singh:2021bwn and Fig. 2 in Singh:2020wsy) and in the LISA triangular space-based interferometer, assuming frequency- and time-independent antenna patterns (Baibhav:2020tma; Marsat:2020rtl). For consistency, we label these modes in Table 3 using the same nomenclature as in Marsat:2020rtl.
These sky modes result from the symmetries in the antenna pattern functions of the three detectors that form ET-$\Delta$. The antenna response functions, which take into account that GW detectors are not equally sensitive to gravitational waves from every direction, can be expressed in the horizontal reference frame, defined by the azimuth ($\beta$) and altitude ($\lambda$) angles. In this frame, the antenna amplitude pattern functions are given by (Regimbau:2012ir):
@@
All these antenna patterns are symmetric under the following transformations: reflection of the sky position across the horizontal plane ($\beta\to-\beta$), rotation about the vertical direction by half a turn ($\lambda\to\lambda+\pi$), and simultaneous rotation of the sky position and polarization angle by a quarter of a turn ($\lambda\to\lambda+\pi/2$, $\psi\to\psi+\pi/2$) (Marsat:2020rtl). Because these transformations produce identical signals in each ET component, they lead to eight degenerate sky positions in the posterior, evenly spaced in $\lambda$ and symmetrically distributed above and below the horizon.
To visualize this effect, we show the total response of each ET component to an incident wave, also called antenna power pattern function: for ET-1, this is given by (Schutz:2011tw):
@@
where the dependency on the polarization angle $\psi$ cancels out. Figure 10 displays the sky location samples in the horizontal reference frame, color-coded by Eq. IV.1. The antenna responses remain identical across all eight modes in all three ET detectors.
The same degeneracies also occur when analyzing the PE of massive black hole binaries (MBHBs) with LISA (Marsat:2020rtl; Pitte:2023ltw; Sun:2025cbo; Yi:2025pxe), provided that the antenna pattern does not change due to Earth’s rotation and the long-wavelength approximation is assumed (Essick:2017wyl). However, in LISA, these approximations can be relaxed, allowing for resolution of these degeneracies (Buscicchio:2021dph). Specifically, an MBHB with a total mass greater than $10^{6}~\mathrm{M}_{\odot}$ can stay in the LISA frequency band for more than 40 hours (Marsat:2020rtl), meaning the rotation of the satellite constellation can have a significant effect.
The long-wavelength approximation (also known as the low-frequency limit) assumes that the antenna pattern functions (Eq. 18 and Eq. 19) act purely as projections of the GW signal onto the detector. This approximation is valid when the length of the detector arms is much shorter than the wavelength of the incoming GW (Misner:1973prb; 1989thyg.book.....H; Rakhmanov:2008is). In reality, this approximation can break down, and the response can also depend on the GW frequency. The frequency at which the long-wavelength approximation becomes important is determined by the unperturbed roundtrip travel time along one arm, also known as the free spectral range (FSR; Marsat:2020rtl; Essick:2017wyl):
@@
Here, $c$ is the speed of light, and for LISA, $L=2.5\times 10^{9}~\mathrm{m}$, which gives $f_{\mathrm{FSR}}=0.06$ Hz. The merger frequency of MBHBs within the LISA frequency band is of the same order of magnitude as $f_{\mathrm{FSR}}$, breaking the long-wavelength approximation (Marsat:2020rtl).
In comparison, for ET-$\Delta$, the free spectral range is $f_{\mathrm{FSR}}=1.5\times 10^{4}~\mathrm{Hz}$, which is at least 3 orders of magnitude higher than the typical merger frequency of the sources considered in this study (see Eq. 2). Moreover, BBH mergers with high detector-frame chirp mass last only a few seconds in the ET-$\Delta$ frequency band. Therefore, we conclude that assuming a time- and frequency-independent antenna pattern is still a valid approximation for high-redshift sources detected with ET-$\Delta$.
IV.2 Geocentric time
Figure 11: Right ascension and declination samples from Dingo-IS converted in azimuth ($\lambda$) and altitude ($\beta$). Color-coded is the geocentric time. See Sec. IV.2 for details.
The sky localization modes discussed in Sec. IV.1 also lead to a bimodal distribution in the geocentric time. This time is calculated by adding a delay to the arrival time at the detector site $(t_{\text{det}})$. This delay depends on the source position in the sky (Singer:2015ema; Harry:2011gre):
$t_{\text{geocent}}=t_{\text{det}}+\frac{\hat{n}\cdot\vec{r}_{\text{ET}}}{c},$ (22)
where $\hat{n}$ is a unit vector pointing to the source in the sky and $\vec{r}_{\text{ET}}$ is the vector from the center of the Earth to the ET-$\Delta$ location.
The scalar product in Eq. 22 depends on the altitude angle $\beta$, defined as $\beta=\arccos(\hat{n}\cdot\hat{r}_{\text{ET}})$, where $\hat{r}_{\text{ET}}$ is the unit vector associated with $\vec{r}_{\text{ET}}$. Of the eight modes described in Sec. IV.1, four are centered around the injected value of $\beta$, and the other four are centered around the opposite angle, $\beta$, as shown in Figs. 10 and 11.
As a result, the geocentric time $t_{\text{geocent}}$ typically shows a bimodal distribution (see Fig. 11). The two modes are most separated when the source is near the local zenith or nadir ($|\beta|\approx\pi/2$). In contrast, when the source is near the local horizon ($\beta\approx 0$), the distribution becomes nearly unimodal (as Event 2 in Appendix B).
IV.3 Higher-order modes and precessing spins
Figure 12: Marginalized one- and two-dimensional posterior distributions for Event 1 with zero noise. We compare waveforms that include aligned spins and higher-order modes (HOM A, dark green) with those that include only the dominant mode ($(2,|2|)$ A, dark magenta), both analyzed using Bilby. Vertical and horizontal lines indicate the true injected values (see Table 2), while the contours represent the 68% and 95% credible regions. Further details are provided in Sec. IV.3.
In the time domain, a GW signal $h=h_{+}+ih_{\times}$ can be expressed as (Newman:1966ub; Berti:2007fi; Kidder:2007rt):